# Po-Heng Chou1, Da-Chih Lin2, Hung-Yu Wei2, Walid Saad3, and Yu Tsao1

### v2

@@ Abstract
In this paper, a measurement-driven framework is proposed for early radio link failure (RLF) prediction in 5G non-standalone (NSA) railway environments. Using 10 Hz metro-train traces with serving and neighbor-cell indicators, we benchmark six models, namely CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet, under varied observation windows and prediction horizons. When the observation window is three seconds, TimesNet attains the highest F1 score with a three-second prediction horizon, while CNN provides a favorable accuracy-latency tradeoff with a two-second horizon, enabling proactive actions such as redundancy and adaptive handovers. The results indicate that deep temporal models can anticipate reliability degradations several seconds in advance using lightweight features available on commercial devices, offering a practical path to early-warning control in 5G-based railway systems.
@@ I Introduction
Despite the rapid deployment of commercial 5G non-standalone (NSA) networks, numerous empirical studies have revealed that their real-world performance often falls short of the theoretical promises of reliability. Measurements across different urban environments show that 5G networks still suffer from non-negligible packet loss and latency fluctuations when compared with mature LTE networks [4, 5]. Such discrepancies are particularly pronounced in high-mobility scenarios, where frequent cell handovers lead to transient disconnections and unstable throughput. Our previous field experiments along the Taipei Metro confirmed this phenomenon: packet losses and excessive latency events predominantly occurred near station areas with intensive handover activities [6]. By analyzing lower-layer signaling messages captured via MobileInsight [7], we identified that handover-related events are the dominant contributors to reliability degradation in 5G NSA networks.
@@
To mitigate these reliability issues, several studies have explored device-side enhancements to improve the stability of 5G connectivity in railway environments. Most of the previous solutions use recovery mechanisms [8, 9] to maintain redundant connections and avoid unnecessary handovers, thereby reducing packet loss and latency. Nevertheless, such methods remain inherently reactive, as they attempt to recover from reliability degradation after it has occurred, without the capability to foresee impending radio link failures (RLFs). In mission-critical railway systems, this lack of early-warning mechanisms poses a persistent risk because communication outages could still occur before mitigation procedures are triggered.
@@
Unlike prior simulation-based or synthetic RLF prediction studies [19, 20, 21], this work presents the first measurement-driven benchmark for early RLF prediction in real 5G NSA metro-train environments. By leveraging field-collected datasets instead of simulated logs, our framework enables quantitative evaluation of learning-based reliability models under authentic mobility, interference, and signaling conditions.
@@
Motivated by these challenges, this paper proposes a measurement-driven, learning-based framework for early failure detection in 5G NSA railway communications. Our objective is to proactively identify potential RLF events before they cause service disruption, enabling preemptive reliability enhancement such as redundancy activation and adaptive handovers.
@@
Based on the practical measurement data collected from metro railway environments, we systematically evaluate the capability of existing learning models to predict reliability degradation in real-world 5G networks. The dataset is sampled at a rate of 10 Hz, corresponding to ten measurement points per second, where each sample includes key radio signal indicators such as the reference signal received power (RSRP) and reference signal received quality (RSRQ) from both serving and neighboring cells. These features capture instantaneous channel conditions as well as short-term fluctuations caused by mobility and handover activities, thereby providing sufficient temporal granularity for modeling reliability dynamics.
In this paper, six representative models are considered, including convolutional neural network (CNN) [10], long short-term memory (LSTM) [11], XGBoost [12], Anomaly Transformer [13], PatchTST [14], and TimesNet [15]. Each model is trained and evaluated under multiple observation windows (1 s, 2 s, 3 s), prediction horizons (1 s, 2 s, 3 s), and sampling schemes with one, two, or three temporal points, either continuous or non-continuous. Rather than proposing a new architecture, we benchmark six representative models under multiple temporal settings, revealing the trade-offs between prediction horizon, observation context, and early-warning reliability in real metro environments.
The main contributions of this paper are summarized as follows:
  • Measurement-driven early warning: We build a supervised framework for proactive RLF prediction using 10 Hz real-world measurements, combining physical-layer indicators (RSRP, RSRQ) with protocol cues to capture reliability dynamics under mobility.
  • Systematic benchmarking: Six representative models (CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, TimesNet) are evaluated under multiple $(T_{s},T_{p})$ settings. TimesNet yields the best overall F1 at $T_{s}=3$ s and $T_{p}=3$ s, while CNN offers a strong accuracy-latency tradeoff at $T_{p}=2$ s.
  • Deployment insights: RLFs account for most packet losses in metro environments and can be predicted up to two seconds ahead, enabling redundancy activation and adaptive handovers in practical 5G railway systems.
@@ II-A Measurement-based Reliability Studies
Empirical studies have revealed that the reliability and latency performance of commercial 5G NSA networks often diverge from the theoretical promises of reliability. Early measurement campaigns showed that, compared with mature LTE systems, 5G networks still exhibit non-negligible packet losses and unstable latency under mobility [4, 5]. To better understand these issues, we conducted extensive field experiments in the Taipei mass rapid transit (MRT) system to analyze the performance of 5G NSA networks under real train operation in our previous work [6]. Our results demonstrated that most packet loss and excessive latency events occur near metro stations, where dense base-station deployment leads to frequent handovers. Statistical analysis further indicated that up to $96\%$ of downlink packet loss rate (PLR) happens during handover-related intervals, confirming that handover is the dominant cause of reliability degradation. These measurement-driven studies established a solid empirical foundation for understanding the reliability bottlenecks of 5G NSA railway communications, motivating the need for predictive mechanisms capable of anticipating RLFs before they occur.
@@ II-B Device-Side Reliability Enhancement
Beyond measurement analysis, several studies have explored device-level techniques to enhance the reliability or throughput of 5G communications in railway scenarios. Most of the previous researches provide a multi-connectivity mechanism [16, 17]. The mechanism enables the UEs to transmit data with multiple replicated paths, thereby reducing the probability of data loss. Those works showed that the proposed method effectively suppresses packet losses and can even achieve zero-loss operation under optimized configurations. Such device-side approaches provide an immediate means to improve network robustness without modifying base-station infrastructure. However, these solutions cannot handle simultaneous link failures caused by the use of the same or highly correlated transmission channels. To ensure safety in mission-critical systems such as railway control, reliability assurance should evolve toward data-driven paradigms that enable early detection of impending RLFs.
@@ II-C Learning-Based Reliability Prediction and Anomaly Detection
In parallel, ML has also been applied to anomaly detection in broader 5G-IoT environments. For instance, [22] introduced a hybrid CNN-LSTM framework that integrates multi-level temporal features to detect traffic anomalies in distributed IoT infrastructures. Such approaches illustrate the general applicability of data-driven reliability modeling across different layers and verticals of 5G systems.
Despite these advances, most existing methods remain validated under simulated or generic network conditions. Their adaptation to mission-critical railway environments, where safety and latency constraints are far more stringent, has not been systematically explored. The present study aims to bridge this gap by establishing a measurement-driven benchmark for early RLF detection in real 5G NSA metro scenarios, enabling quantitative assessment of existing models under realistic mobility and signal fluctuation patterns.
@@ III-A Experimental Environment and Tools
In each experiment, the UE was implemented using commercial smartphones connected via universal serial bus (USB) to a laptop, which served as the data client. A remote server located at National Taiwan University acted as the communication endpoint. The downlink (DL) and uplink (UL) data streams were transmitted using the user datagram protocol (UDP), with a fixed payload size of 250 bytes and a transmission rate of 200 kbps. This configuration provided a lightweight yet continuous traffic load suitable for reliability evaluation while minimizing protocol overhead.
@@
For control-plane observation, we employed MobileInsight [7], an open-source tool that extracts LTE and 5G NR signaling events from modem logs. This allowed us to record detailed radio resource control (RRC) procedures, such as connection setup, reconfiguration, and handover events. The integration of transport-layer packet traces with low-layer signaling messages enabled a comprehensive view of how physical and control-layer behaviors jointly influence communication reliability in high-mobility 5G NSA networks.
@@ III-B Observations
Fig. 1 summarizes the distribution of downlink packet loss across various event categories observed in our experiments. The analysis indicates that two specific events, NAS Recovery (NASR) and Medium Access Control Configuration Failure (MCGF), exhibit the highest packet loss per event. When the overall frequency and impact are combined, the five most influential events are MCGF, Master Node Base Handover (MNBH), Secondary Cell Group Modification (SCGM), End Node Base Handover (ENBH), and NASR, in descending order of contribution.
A critical pattern emerges when correlating these events with radio link status: both MCGF and NASR are classified as RLF events. Together, they account for 58.3% of all downlink packet losses, with MCGF alone responsible for over half of the total (54.6%). This finding confirms that RLFs are the dominant source of reliability degradation in 5G NSA railway communications.
These results suggest that device-side mechanisms can mitigate ongoing failures, yet most losses originate from abrupt and unpredictable RLF events. Identifying such events in their early stages would allow controllers or user equipment to trigger redundancy and adaptive procedures before a full breakdown. Developing learning-based early warning is therefore essential for improving the reliability and resilience of 5G railway communications.
@@ IV RLF Prediction Design
To proactively address radio link failure (RLF) events, we formulate the task as a supervised time-series classification problem. As illustrated in Fig. 2, at each time point $t$ (denoted as Now), the predictor takes a sliding observation window of length $T_{s}$ seconds as input, consisting of historical measurements sampled every $I_{s}$ seconds. The model then outputs the probability that an RLF will occur within the subsequent prediction horizon of $T_{p}$ seconds. For binary classification, a sample is labeled $y=1$ if any RLF occurs within $(t,\,t+T_{p}]$, and $y=0$ otherwise. For the multi-interval variant, a one-hot vector indicates the sub-interval of $(t,\,t+T_{p}]$ in which the next RLF appears. RLF samples form the positive class.
@@
RLF events are inherently rare in typical operational settings. Our experiments, conducted with a sampling interval of $I_{s}=0.1$ s (10 Hz), show an approximate ratio of one RLF sample per 500 non-RLF samples, resulting in a highly imbalanced dataset that complicates accurate prediction. This fine-grained sampling rate was intentionally chosen because network performance metrics can fluctuate at millisecond timescales; a 0.1-second interval allows us to capture these subtle variations more effectively, thereby improving the model’s ability to identify early precursors of RLF events. To address this challenge, we examined several widely used techniques for dealing with imbalanced data, including synthetic minority over-sampling (SMOTE), random downsampling to achieve a target ratio of 30:1 (non-RLF to RLF), and class-weighted loss functions that emphasize the minority class during training.
@@ V-A Implementation Environment and Model Setup
All experiments were implemented in Python 3.10 using TensorFlow and XGBoost libraries within Visual Studio Code. Model training and evaluation were executed on a 13th Gen Intel® Core™ i5-13400 CPU (10 cores, 16 threads, base 2.5 GHz, boost 4.6 GHz, 12 MB cache, 65 W TDP) without GPU acceleration. Six representative models, including CNN, LSTM, XGBoost, TimesNet, Anomaly Transformer, and PatchTST, were benchmarked under identical training settings for fair comparison. All models adopted the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, batch size of 128, and early stopping based on the validation Area Under Curve (AUC) (patience =10=10). A learning rate reduction strategy (ReduceLROnPlateau) with a minimum learning rate of $1\times 10^{-5}$ and class weighting for imbalance correction were applied. Training was conducted for up to 60 epochs. Detailed layer structures, preprocessing scripts, and hyperparameter configurations are available in the project repository. [^1] Source code and configuration files are available at: https://github.com/BoneZhou/RLF-Prediction.
@@ V-B Evaluation Setup
Unless otherwise stated, we report accuracy, precision, recall, and F1 at the F1-optimal decision threshold selected on the validation set by sweeping $\tau\in\{0.1,0.2,\ldots,0.9\}$. RLF is treated as the positive class. We consider observation windows $T_{s}\in\{1,2,3\}$ s and prediction horizons $T_{p}\in\{1,2,3\}$ s with a sampling rate of $I_{s}=0.1$ s (10 Hz).

### v4

# Po-Heng Chou1,3, Da-Chih Lin2, Hung-Yu Wei2, Walid Saad3, and Yu Tsao1

@@ Abstract
This paper presents a measurement-driven study of early warning for reliability breakdown events in 5G non-standalone (NSA) railway networks. Using 10 Hz metro-train measurement traces with serving- and neighbor-cell indicators, we benchmark six representative learning models, including CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet, under multiple observation windows and prediction horizons. Rather than proposing a new prediction architecture, this study develops a measurement-driven benchmark to quantify the feasibility and operating trade-offs of seconds-ahead reliability prediction in 5G NSA railway environments. Experimental results show that learning models can anticipate RLF-related reliability breakdown events seconds in advance using lightweight radio features available on commercial devices. The presented benchmark provides insights for sensing-assisted communication control and offers an empirical foundation for integrating sensing and analytics into future mobility control.
@@ I Introduction
Despite the rapid deployment of commercial 5G non-standalone (NSA) networks, numerous empirical studies have revealed that their real-world performance often falls short of the theoretical promises of reliability. Measurements across different urban environments show that 5G networks still suffer from non-negligible packet loss and latency fluctuations when compared with mature LTE networks [4]. Such discrepancies are particularly pronounced in high-mobility scenarios, where frequent cell handovers lead to transient disconnections and unstable throughput. Our previous field experiments along the Taipei Metro confirmed this phenomenon: packet losses and excessive latency events predominantly occurred near station areas with intensive handover activities [5]. By analyzing lower-layer signaling messages captured via MobileInsight [6], we identified that handover-related events are the dominant contributors to reliability degradation in 5G NSA networks. From a 6G viewpoint, early RLF warning can be regarded as device-side network sensing that supports sensing-assisted mobility control. In this context, mobile devices can be viewed as distributed sensing nodes that continuously monitor network reliability under high-mobility conditions.
@@
To mitigate these reliability issues, several studies have explored device-side enhancements to improve the stability of 5G connectivity in railway environments. Most prior solutions rely on recovery mechanisms [7, 8] to maintain redundant connections and avoid unnecessary handovers, thereby reducing packet loss and latency. While such operational studies provide important insights into failure handling and recovery in commercial 5G deployments, they remain fundamentally reactive because mitigation is triggered only after degradation has emerged. In contrast, this work focuses on the complementary problem of seconds-ahead early warning using lightweight radio features on commercial devices. The goal of this paper is therefore not to replace failure recovery mechanisms, but to quantify the feasibility and operating trade-offs of predictive warning before reliability breakdown occurs.
@@
Unlike prior simulation-based or synthetic RLF prediction studies [18, 19, 20], this work presents a measurement-driven benchmarking study for early prediction of RLF-related reliability breakdown events in a real 5G NSA metro-train environment. By leveraging field-collected datasets instead of simulated logs, the proposed framework enables a quantitative evaluation of learning-based reliability models under authentic mobility, interference, and signaling conditions.
@@
Motivated by these challenges, this paper presents a measurement-driven framework for early detection of reliability breakdown events in 5G NSA railway communications. Our objective is to identify impending reliability breakdown events before they cause observable service disruption, thereby supporting preemptive reliability enhancement such as warning generation, redundancy preparation, and adaptive handover assistance.
@@
Based on the practical measurement data collected from metro railway environments, we systematically evaluate the capability of existing learning models to predict reliability degradation in real-world 5G networks. The dataset is sampled at a rate of 10 Hz, corresponding to ten measurement points per second, where each sample includes key radio signal indicators such as the reference signal received power (RSRP) and reference signal received quality (RSRQ) from both serving and neighboring cells. These features capture instantaneous channel conditions as well as short-term fluctuations caused by mobility and handover activities, thereby providing sufficient temporal granularity for modeling reliability dynamics. We benchmark six representative models, namely CNN [9], LSTM [10], XGBoost [11], Anomaly Transformer [12], PatchTST [13], and TimesNet [14], under multiple observation windows and prediction horizons, where $T_{s}$ and $T_{p}$ denote the observation window and prediction horizon, respectively.
The key contributions of this study are summarized as follows:
  • Measurement-driven early-warning benchmark based on device-side sensing: We present a supervised early-warning framework using 10 Hz real-world railway measurements collected from a commercial 5G NSA metro environment, leveraging lightweight radio indicators (RSRP, RSRQ) and neighbor-cell dynamics.
  • Systematic model comparison under temporal design choices: Six representative models (CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet) are evaluated under multiple $(T_{s},T_{p})$ settings to characterize the trade-offs among sensing context, prediction horizon, and alarm reliability.
  • Operational insights for proactive mobility control: We show that RLF-related reliability breakdown events can be anticipated seconds ahead, and we analyze deployment-relevant trigger policies that can support proactive actions such as redundancy activation and adaptive handovers.
@@ II-A Measurement-based Reliability Studies
Empirical studies have revealed that the reliability and latency performance of commercial 5G NSA networks often diverge from the theoretical promises of reliability. Early measurement campaigns showed that, compared with mature LTE systems, 5G networks still exhibit non-negligible packet losses and unstable latency under mobility [4]. To better understand these issues, we conducted extensive field experiments in the Taipei mass rapid transit (MRT) system to analyze the performance of 5G NSA networks under real train operation in our previous work [5]. Our results demonstrated that most packet loss and excessive latency events occur near metro stations, where dense base-station deployment leads to frequent handovers. Statistical analysis further indicated that up to $96\%$ of downlink packet loss rate happens during handover-related intervals, confirming that handover is the dominant cause of reliability degradation. These measurement-driven studies established a solid empirical foundation for understanding the reliability bottlenecks of 5G NSA railway communications, motivating the need for predictive mechanisms capable of anticipating RLFs before they occur.
@@ II-B Device-Side Reliability Enhancement
Beyond measurement analysis, several studies have explored device-level techniques to enhance the reliability or throughput of 5G communications in railway scenarios. Representative approaches include multi-connectivity and recovery-oriented mechanisms [15, 16, 7, 8], which maintain redundant connections, avoid unnecessary handovers, or recover from degraded links after failure symptoms emerge. These methods provide an immediate means to improve network robustness without modifying base-station infrastructure. However, they are still fundamentally reactive and may not prevent abrupt reliability breakdown before mitigation is triggered. This motivates data-driven early-warning mechanisms that can complement recovery-oriented designs by predicting impending failures in advance.
@@ II-C Learning-Based Reliability Prediction and Anomaly Detection
In parallel, ML has also been applied to anomaly detection in broader 5G-IoT environments. For instance, [21] introduced a hybrid CNN-LSTM framework that integrates multi-level temporal features to detect traffic anomalies in distributed IoT infrastructures. Such approaches illustrate the general applicability of data-driven reliability modeling across different layers and verticals of 5G systems. Overall, these studies motivate a measurement-grounded view of early RLF warning as network sensing and analytics that can support sensing-assisted mobility control in future integrated communication systems.
@@ III-A Experimental Environment and Tools
In each experiment, the UE was implemented using commercial smartphones connected via universal serial bus (USB) to a laptop, which served as the data client. A remote server located at National Taiwan University acted as the communication endpoint. The downlink and uplink data streams were transmitted using the user datagram protocol (UDP), with a fixed payload size of 250 bytes and a transmission rate of 200 kbps. This configuration provided a lightweight yet continuous traffic load suitable for reliability evaluation while minimizing protocol overhead.
@@
For control-plane observation, we employed MobileInsight [6], an open-source tool that extracts LTE and 5G NR signaling events from modem logs. This allowed us to record detailed radio resource control (RRC) procedures, such as connection setup, reconfiguration, and handover events. The integration of transport-layer packet traces with low-layer signaling messages enabled a comprehensive view of how physical and control-layer behaviors jointly influence communication reliability in high-mobility 5G NSA networks. Although the campaign was conducted on a single metro line and operator setup, the route spans urban, semi-open, and underground segments, providing heterogeneous propagation conditions for an initial measurement-driven benchmark.
@@ III-B Observations
Fig. 1 summarizes the distribution of downlink packet loss across various event categories observed in our experiments. The analysis indicates that two specific events, Non-Access Stratum Recovery (NASR) and Master Cell Group Failure (MCGF), exhibit the highest packet loss per event. When the overall frequency and impact are combined, the five most influential events are MCGF, Master eNB Handover (MNBH), Secondary Cell Group Modification (SCGM), eNB Handover (ENBH), and NASR, in descending order of contribution.
A critical pattern emerges when correlating these events with radio link reliability: both MCGF and NASR correspond to reliability breakdown events that are strongly associated with radio link failure behavior in the measured 5G NSA network. Together, they account for 58.3% of all downlink packet losses, with MCGF alone responsible for over half of the total (54.6%). This finding confirms that RLF-related reliability breakdown events are the dominant source of reliability degradation in the measured 5G NSA railway environment.
These observations indicate that many reliability breakdown events are preceded by measurable radio-layer fluctuations, suggesting that lightweight device-side indicators can support early warning before a full breakdown. Developing learning-based early warning is therefore essential for improving the reliability and resilience of 5G railway communications.
@@ IV RLF Prediction Design
In this paper, the positive class is defined using stack-level events observed from operational logs, primarily MCGF and NASR. These events do not strictly correspond to the formal 3GPP RLF definition based on out-of-sync indications and timer expiry. However, they represent practically observable reliability breakdown events that lead to packet loss or service interruption in commercial 5G NSA operation. Therefore, the prediction target in this work is more precisely interpreted as RLF-related reliability breakdown events. To proactively address RLF events, we formulate the task as a supervised time-series classification problem. As illustrated in Fig. 2, at each time point $t$ (denoted as Now), the predictor takes a sliding observation window of length $T_{s}$ seconds as input, consisting of historical measurements sampled every $I_{s}$ seconds. The model then outputs the probability that an RLF-related reliability breakdown event will occur within the subsequent prediction horizon of $T_{p}$ seconds. For binary classification, a sample is labeled $y=1$ if any target reliability breakdown event occurs within $(t,\,t+T_{p}]$, and $y=0$ otherwise. For the multi-interval variant, a one-hot vector indicates the sub-interval of $(t,\,t+T_{p}]$ in which the next RLF appears. Positive samples correspond to RLF-related reliability breakdown events.
@@
RLF events are inherently rare in typical operational settings. Our experiments, conducted with a sampling interval of $I_{s}=0.1$ s (10 Hz), show an approximate ratio of one RLF sample per 500 non-RLF samples, resulting in a highly imbalanced dataset that complicates accurate prediction. This fine-grained sampling rate was intentionally chosen because network performance metrics can fluctuate at millisecond timescales; a 0.1-second interval allows us to capture these subtle variations more effectively, thereby improving the model’s ability to identify early precursors of RLF events. To mitigate this challenge, we considered several widely used strategies for imbalanced learning, including synthetic minority over-sampling (SMOTE), random downsampling to achieve a target ratio of 30:1 (non-RLF to RLF), and class-weighted loss functions that emphasize the minority class during training.
@@ V Performance Results
(a) CNN [9] with $T_{p}=2$ s
(b) XGBoost [11] with $T_{p}=2$ s
(c) TimesNet [14] with $T_{p}=3$ s
Figure 3: Prediction hits of the top-3 models based on Tables I. Profiles use $T_{s}=3$ s and $I_{s}=0.1$ s; CNN and XGBoost are shown with $T_{p}=2$ s, and TimesNet with $T_{p}=3$ s. Red markers indicate correct predictions, i.e., time slots where an RLF-related reliability breakdown event was predicted to occur within the next $T_{p}$ seconds.
@@ V-A Implementation Environment and Model Setup
All experiments were implemented in Python 3.10 using TensorFlow and XGBoost libraries within Visual Studio Code. Model training and evaluation were executed on a 13th Gen Intel® Core™ i5-13400 CPU (10 cores, 16 threads, base 2.5 GHz, boost 4.6 GHz, 12 MB cache, 65 W TDP) without GPU acceleration. Six representative models, including CNN, LSTM, XGBoost, TimesNet, Anomaly Transformer, and PatchTST, were benchmarked under identical training settings for fair comparison. All models adopted the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, batch size of 128, and early stopping based on the validation Area Under Curve (AUC) (patience =10=10). A learning rate reduction strategy (ReduceLROnPlateau) with a minimum learning rate of $1\times 10^{-5}$ and class weighting for imbalance correction were applied. Training was conducted for up to 60 epochs. Detailed layer structures, preprocessing scripts, and hyperparameter configurations are available in the project repository. [^1] Source code and configuration files are available at: github.com/BoneZhou/RLF-Prediction.
@@ V-B Evaluation Setup
Unless otherwise stated, we report accuracy, precision, recall, and F1 at the F1-optimal decision threshold selected on the validation set by sweeping $\tau\in\{0.1,0.2,\ldots,0.9\}$. The positive class corresponds to RLF-related events. We consider observation windows $T_{s}\in\{1,2,3\}$ s and prediction horizons $T_{p}\in\{1,2,3\}$ s with a sampling rate of $I_{s}=0.1$ s (10 Hz).
Because the dataset is highly imbalanced (approximately 1:500 positive-to-negative ratio), accuracy and ROC-oriented metrics alone may overestimate practical utility. Therefore, in addition to accuracy, precision, recall, and F1, we also evaluate deployment-oriented alarm reliability using PR-AUC, false alarms per minute, and recall under fixed false-alarm budgets whenever applicable. These metrics better reflect the operational usefulness of early-warning decisions in highly imbalanced railway communication settings.
Beyond sample-level classification scores, an early-warning system should also be assessed from an operational perspective. In particular, excessive false alarms may trigger unnecessary redundancy activation, handover attempts, or control signaling overhead in safety-critical railway systems. Therefore, we additionally consider alarm-centric metrics such as PR-AUC, false alarms per minute, and event-level hit rate under different trigger policies.